OpenStudy (anonymous):

Solve for x: |2x + 6| - 4 = 20 a. x = 9 and x = 11 b. x = -9 and x = 15 c. x = 9 and x = - 15 d. no solution okay I got x = 9 but i dont know what the other number is?

jimthompson5910 (jim_thompson5910):

you got |2x + 6| = 24 right? then you got 2x + 6 = 24 ---> x = 9

jimthompson5910 (jim_thompson5910):

since |2x + 6| = 24, it breaks down into these two equations 2x + 6 = 24 OR 2x + 6 = -24

jimthompson5910 (jim_thompson5910):

because |x| = k breaks down into x = k or x = -k (k is positive) example: |x| = 7 ---> x = 7 or x = -7